# LoRA-Mixer: Coordinate Modular LoRA Experts Through Serial Attention Routing

### v1

@@ 4.1 Experimental Setup
Datasets. To evaluate LoRA-Mixer, we selected seven publicly available benchmarks. For medical question answering, we used the Medical-QA dataset. For commonsense reasoning, we adopted ARC-E [30] and ARC-C [30], both of which focus on multiple-choice questions. For natural language understanding, we used SST2 [31] and CoLA [31], where SST2 is used for sentiment classification and CoLA is used for grammatical judgment. For mathematical reasoning, we chose the GSM8K [29] dataset, which contains thousands of elementary school math problems that require multi-step solutions. Finally, to evaluate encoding ability, we chose the HumanEva [19]l dataset.
Baselines. We chose three open source LLMs - LLaMA3-8B [26], Mistral-7B [27] and Falcon-Mamba-7B [28]. LLaMA3-8B and Mistral-7B both rely on the Transformer backbone network; Falcon-Mamba-7B is built entirely on the Mamba paradigm. In addition, we compared LoRA-Mixer with other state-of-the-art methods, including MoLE [14],MixLoRA [13],LoraHub [15],LoRA-LEGO [16].
Evaluation metrics We use three evaluation metrics to measure the performance. Specifically, for GSM8K [29], ARC-E [30], ARC-C [30], SST2 [31], and CoLA [31], we use ACC to measure performance. For the HumanEval dataset [19], we adopt the $Pass@1$ metric, which represents the ability of a single generated answer to correctly solve the task. Finally, considering the domain-specific freedom and rigor required by the Medical-QA dataset, we use DeepSeek-R1 [22] to evaluate completeness, correctness, and logical clarity, and report the final percentage scores.
@@ 4.2 Comparisons
Table1 shows the performance indicators of the three basemodels on various tasks. To ensure the reliability of the experimental results, we ran all experiments three times and took the average value.
@@
Falcon-Mamba-7B[28] 73.67 82.42 92.81 52.54 77.61 68.78 29.29
Mistral-7B[27] 66.32 71.21 85.24 40.83 80.00 61.50 27.95
LLaMA3-8B[26] 78.47 79.14 93.12 57.92 88.45 78.65 52.44
We compare our method with the state-of-the-art methods [14, 13, 15, 16] on seven datasets, and the experimental results are shown in Table 2. As can be seen from Table 2, our method achieves better performance than the baseline on most tasks. For the Falcon-Mamba , our method significantly outperforms other baselines on all tasks. For model details, please refer to Appendix B.
Table 2: Comparison of our LoRA-Mixer with LoRAHub [15], MoLE [14], and Mix-LoRA [13] across seven tasks (best scores in bold). Note that MixLoRA is excluded from Falcon-Mamba due to its Transformer-specific design.
@@
LoRAHub [15] 69.17 75.73 90.21 44.94 81.14 69.21 32.60
MOLE [14] 71.07 78.51 94.17 45.31 85.68 68.77 35.37
MixLoRA [13] 69.74 78.61 93.44 45.50 85.42 69.15 33.80
@@
LoRAHub [15] 78.11 79.84 92.77 59.10 87.13 80.14 52.83
MOLE [14] 78.43 81.37 94.18 63.81 88.15 81.77 55.87
MixLoRA [13] 79.87 80.67 94.22 64.44 88.70 82.90 55.49
@@
Table 3: Evaluation of LoRA-Mixer on LoRAs sourced from Internet on five GLUE tasks, the base model is Flan-T5[36].
@@
Table 4: Comparison of LoRA-Mixer and LoRA-LEGO. Results for LoRA-LEGO are from its paper [16].
@@
Since LoRA-LEGO [16] is not open-sourced, to compare fairly, we use LLaMA2-7B [33] as the base model and conduct experiments on four tasks including CoLA, SST2, MRPC and RTE from LoRA-LEGO paper. Lora’s configuration employs a low rank of $r=6$ and a scaling factor of $\alpha=12$. The experimental results are shown in Table 4. From the results, it can be seen that our method outperforms LoRA-LEGO in three of the four tasks.
Considering that Mistral-7B and LLaMA3-8B have exactly the same model architecture, we directly migrate the parameters trained on Mistral-7B to LLaMA3-8B without any fine-tuning and adaptation, and conduct experiments on three datasets: ARC-E, ARC-C, and GSM8K. The results are shown in Table 5 . It is worth noting that we use the Zero-Shot CoT method to test the base model in the GSM8K task, and the results under different Few-Shot settings are also shown in Table 5.
@@
Interestingly, we observed that we achieved better performance than the LLaMA3-8B baseline on two of the three tasks. This cross-model migration verifies that the LoRA expert and the learned routing function are not tightly coupled with a specific base model, making it possible to share experts between models with the same architecture.
@@ 4.3 Testing on LoRAs Sourced from Internet
To verify flexibility and plugin and play nature of LoRA-Mixer, we test our LoRA-Mixer on LoRAs sourced from Internet. Specifically, we download five distinct LoRAs from LoRAHub [15], which were trained on SST2, CoLA, MRPC, RTE, and QQP, respectively (please refer to Appendix E for more details). We use the Flan-T5 [36] model as a base model and mount the LoRAs without any modification, and collect $2K$ mixed data for routing training. The results are shown in Table 4. LoRA-Mixer achieved better performance on the four tasks, confirming that LoRA-Mixer’s potential for product-ready multitask applications.
@@ 4.4 Ablation Study
Impact of LoRA Rank. To evaluate the impact of rank in low-rank adaptation, we conducted experiments on $r=16$, $r=32$, $r=64$, and $r=128$, while keeping all other hyperparameters (e.g., dropout rate, learning rate) unchanged. The results of $r=64$ can be found in Table 2. We place the remaining results in Appendix A.
@@
Expert Load Analysis. To analyze the overall load of experts, we uniformly sampled 1K data for seven benchmarks, including Medical, CoLA, SST2, GSM8K, ARC-E, ARC-C, and Humaneval. We report the average load of each expert on these $1K$ data, as shown in Figure 4. The activation rates of different experts are quite balanced, ranging from 15% - 18%, but in different tasks, the expert load reflects a kind of "perception" ability, and the expert load of specific tasks is higher than that of other experts, as shown in Figure  5. This shows that our routing mechanism effectively avoids expert collapse and achieves a balance of expert utilization between different tasks.
The impact of Top-$K$. To explore the effect of K values on Top-K routing, we used Falcon-Mamba as the basemodel to experiment on SST-2 and CoLA. The experimental results are shown in Figure  3.
Figure 3: Top-K Routing Impact.
Figure 4: Expert Assignment Overview.
As $K$ increases from 1 to 3, we observe that the accuracy of both tasks improves, indicating that using multiple experts allows the model to obtain complementary information. However, further increasing the value of K does not guarantee better results, but may degrade the performance. Therefore, the setting of $K$ is crucial for the MoE model. How to set or dynamically learn the most appropriate $K$ value is a direction worthy of further research in the future.
Table 6: The impact of LoRA-Mixer optimization on individual LoRAs (LoRA means adding independently trained LoRA, Expert means LoRA optimized by LoRA-Mixer).
Task LoRA w/o Expert w/ Expert Gap (Expert)
SST-2 95.30 94.70 95.41 +0.71
CoLA 81.50 80.15 82.22 +2.07
GSM8K 65.14 60.17 65.53 +5.36
HumanEval 55.61 53.39 57.32 +3.93
Enhanced Expert Performance Analysis. In order to verify the enhanced performances of individual expert after LoRA-Mixer optimization. We selected four tasks, GSM8K, SST2, CoLA and HumanEval, for experiments. The results are shown in Table 6. We can find that LoRAs optimized by LoRA-mixer exhibit improved performance, especially in the GSM8K task, where the performance improved by 5.36% after adding mathematical experts. This result confirms the enhanced ability of each individual LoRA expert after LoRA-Mixer optimization.
Cross-domain QA To evaluate the cross-domain generalization ability of LoRA-Mixer, we constructed two question-answering datasets: Medical-Mathematics and Mathematics-Coding. Each dataset contains 200 samples generated by DeepSeek-R1. These questions are more challenging. In the Medical-Mathematics dataset, the model needs to provide effective medical advice and corresponding calculations. In the Mathematics-Coding dataset, the model needs to generate correct Python code based on mathematical problems. The evaluation results of DeepSeek-R1 are shown in Table  7.
Table 7: Cross-domain performance of LoRA-Mixer on LLaMA3-8B [26].
Task Base LoRAHub [15] MOLE [14] MixLoRA [13] LoRA-Mixer
Math-Medical 69.88 70.53 72.11 72.74 73.41
Math-Coding 59.37 61.08 62.24 63.10 63.46
The impact of the RSL. We study the impact of our proposed RSL loss function on the LoRA-Mixer framework. RSL has two major advantages.
First, it enables routers to achieve global load balancing while maintaining strong input perception, which is a key factor to fully exploit the potential of sparse expert models. Second, compared with traditional load balancing loss functions, RSL significantly reduces the amount of training data required for effective router optimization.
@@
To verify the first conclusion, we conducted experiments on three tasks: Medical, GSM8K, and HumanEval. As shown in Figure 5, after using RSL loss, LoRA-Mixer always assigns higher activation weights to experts related to the target task, which reflects the router’s strong domain perception and adaptive specialization capabilities. In contrast, using only auxiliary loss, the router will evenly distribute experts regardless of the semantics of the input content, which will result in the potential of the relevant experts not being fully developed, causing a performance bottleneck.
@@
As shown in Table 4.4, LoRA-Mixer achieves comparable performance using only 51.62% of the routing supervision data required by traditional auxiliary loss methods. We attribute this data efficiency to the dual regularization mechanism introduced by RSL. Specifically, a global consistency term $\bar{\mathbf{p}}^{\top}\bar{\mathbf{f}}$ aligns the expected routing probabilities with actual expert utilization, while a local token-level entropy penalty encourages diverse and selective expert activation. This synergistic design mitigates the overly uniform expert usage common in auxiliary loss, promoting both expert specialization and routing sparsity. As a result, the model maintains robust and adaptive expert assignment even under limited supervision. A detailed theoretical justification is provided in Appendix C.
5 Conclusion and Discussion

### v2

@@ 4.1 Experimental Setup
Datasets We tested RSL and LoRA-Mixer on 15 public benchmarks: MedicalQA, ARC-E, ARC-C, GSM8K, GLUE, PIQA, BoolQ, HumanEval, and HellaSwag. These datasets cover five different domains. For detailed dataset information, please refer Appendix 11.
Baseline We selected LLaMA3-8B, Mistral-7B, and Falcon-Mamba-7B as the basemodel. LLaMA3-8B and Mistral-7B are Transformer architectures, while Falcon-Mamba is a pure SSM architecture. We compare LoRA-Mixer and RSL with other state-of-the-art methods, including MoLE Wu et al. (2024), MixLoRA Li et al. (2024a), LoraHub Huang et al. (2023), LoRA-LEGO Zhao et al. (2024) and PHATGOOSE Muqeeth et al. (2024). We also select strong baselines that specifically optimize auxiliary losses to demonstrate the effectiveness of RSL, including GMoE Bai et al. (2024), AESL Guo et al. (2025b), and DsMoE Pan et al. (2024). It is worth noting that MoLE and MixLoRA also optimize auxiliary losses. For parameter, training and inference analysis, please refer to A.4 A.7.
Metrics For HumanEval, we use the $Pass@1$ metric. Considering the domain-specific freedom and rigor required by the Medical-QA dataset, we use DeepSeek-R1 for evaluation. For the remaining tasks, we use ACC as the metric. To reduce randomness, all experiments are run three times and the average reported.
@@ 4.2 Comparisons
Falcon-Mamba-7B 73.67 82.42 92.81 52.54 77.61 68.78 29.29
Mistral-7B 66.32 71.21 85.24 40.83 80.00 61.50 27.95
LLaMA3-8B 78.47 79.14 93.12 57.92 88.45 78.65 52.44
Table1 shows the performance indicators of the three basemodels on various tasks. We compare our approach with the state-of-the-art methods on seven datasets. The experimental results are shown in Table 2. Our approach outperforms the baselines on most tasks. For the Falcon-Mamba dataset, our approach significantly outperforms the baselines on all tasks. For model details, see Appendix A.11.
We use the LLaMA2-7B from the LoRA-LEGO paper as the basemodel and conduct experiments on the four tasks: CoLA, SST2, MRPC, and RTE. The LoRA configuration uses $r=6$ and $\alpha=12$. The experimental results are shown in Table 4. As can be seen, our method outperforms LoRA-LEGO on three of the four tasks.
Table 2: Comparison of our LoRA-Mixer with LoRAHub, MoLE, and Mix-LoRA across seven tasks (best scores in bold). Note that MixLoRA is excluded from Falcon-Mamba due to its Transformer-specific design.
@@
LoRAHub 69.17 75.73 90.21 44.94 81.14 69.21 32.60
MOLE 71.07 78.51 94.17 45.31 85.68 68.77 35.37
MixLoRA 69.74 78.61 93.44 45.50 85.42 69.15 33.80
@@
LoRAHub 78.11 79.84 92.77 59.10 87.13 80.14 52.83
MOLE 78.43 81.37 94.18 63.81 88.15 81.77 55.87
MixLoRA 79.87 80.67 94.22 64.44 88.70 82.90 55.49
@@
Table 3: Evaluation of LoRA-Mixer on LoRAs sourced from Internet on five GLUE tasks, the base model is Flan-T5Chung et al. (2024).
@@
Table 4: Comparison of LoRA-Mixer and LoRA-LEGO. Results for LoRA-LEGO are from its paper.
@@
Considering that Mistral-7B and LLaMA3-8B have the same architecture, we directly migrate the parameters trained on Mistral-7B to LLaMA3-8B without any fine-tuning and adaptation, and conduct experiments on three datasets: ARC-E, ARC-C, and GSM8K. The results are shown in Table 5 . It is worth noting that we use the Zero-Shot CoT method to test the basemodel in the GSM8K task, and the results under different Few-Shot settings are also shown in Table 5.
@@
Interestingly, we observe that we outperform the LLaMA3-8B on two of the three tasks. This cross-model transfer validates the design motivation of LoRA-Mixer and demonstrates that the routing learned via RSL is extremely robust and transferable, making it possible to share experts between models with the same architecture.
To further demonstrate the performance of LoRA-Mixer, we conducted experiments on BoolQ, HellaSwag, and PIQA using LLaMA3-8B, with $r=32$ for LoRA. The results are shown in Table 7. It can be seen that LoRA-Mixer can effectively coordinate the capabilities of LoRAs to enhance the basemodel. To test the generalization of LoRA-Mixer, we conducted experiments with PHATGOOSE on 9 datasets, six of which were in-distribution experiments and three were out-of-distribution experiments. Here we only show the OOD experimental results because OOD tasks can reflect the generalization ability of LoRA-Mixer. For the results of the in-distribution experiment, please refer to A.14. The results show that LoRA-Mixer has excellent generalization ability, which means that through RSL, routing can make professional decisions and achieve better performance.
Table 6: OOD comparison across datasets for Phatgoose, and LoRA-Mixer.
Dataset Base Phatgoose LoRA-Mixer
QQP 68.84 69.28 69.47
RTE 70.88 69.87 71.31
MRPC 74.73 74.59 74.93
Table 7: Results on BoolQ, HellaSwag, and PIQA. Base is LLaMA3-8B.
Dataset Base LoRA LoRA-Mixer Gap
BoolQ 71.25 74.46 79.37 +4.91
HellaSwag 75.33 77.39 82.41 +5.02
PIQA 78.47 80.71 84.94 +4.23
@@ 4.3 Testing on LoRAs Sourced from Internet
To verify that LoRA-Mixer can reuse LoRAs with minimal effort, we tested it on LoRAs from the internet. We downloaded five different LoRAs from LoRAHub Huang et al. (2023) and trained them on SST2, CoLA, MRPC, RTE, and QQP (see Appendix A.15 for details). We used the Flan-T5 model as the base model, froze the LoRAs parameters, and collected an additional 2k mixed data points for routing training, which is independent of the LoRAs training data. The results are shown in Table 4. LoRA-Mixer achieved excellent performance on all four tasks, demonstrating its potential for production-grade multi-task applications.
4.4 Comparison with other optimized routing losses.
Although we have compared RSL with other optimized routing losses (MoLE, MixLoRA), to further demonstrate the effectiveness of RSL, we choose three strong baselines for comparison: GMoE, Ds-MoE, and AESL, all of which are specifically optimized for routing losses. It is worth noting that all experiments are conducted with the same training data (2k), and the only difference is the routing loss. As shown in Table 4.4, RSL significantly outperforms other strong baselines in low-data-resource scenarios.
Table 8: Comparison with GMoE, DS-MoE, and AESL under the same training data (2k) and LoRAs parameters. Demonstrates the advantages of RSL under low data resources.
Task GMoE DS-MoE AESL RSL
SST-2 91.38 92.45 92.64 95.41
CoLA 79.57 79.83 80.42 82.22
ARC-E 85.65 85.32 86.24 89.88
ARC-C 76.42 78.45 79.88 83.24
HumanEval 46.37 48.92 50.46 57.32
Figure 3: Expert Assignment Overview.
@@ 4.5 Ablation Study
Impact of LoRA Rank. To evaluate the impact of $r$ in LoRAs, we conducted experiments on $r=16$, $r=32$, $r=64$, and $r=128$, while keeping all other hyperparameters (e.g., dropout rate, learning rate) unchanged. The results of $r=64$ can be found in Table 2. We place the remaining results in Appendix A.12.
@@
Expert Load Analysis. To analyze the overall load, we uniformly sampled 1K data from seven benchmarks (MedicalQA, CoLA, SST2, GSM8K, ARC-E, ARC-C, and Humaneval). We report the average load per expert on this 1K data, as shown in Figure 3 . The activation rates of different experts are relatively balanced, ranging from 15% to 18%. However, across tasks, expert load reflects a "perception" ability, with experts on certain tasks having higher loads than others, as shown in Figure  4 . This demonstrates that our routing mechanism effectively avoids expert collapse and achieves balanced expert utilization across different tasks.
The impact of K and Cross-domain QA. We also analyzed the impact of K in Top-K. In general, increasing K within an appropriate range can improve model performance. For details, please refer to A.3. Furthermore, to evaluate LoRA-Mixer’s generalization capabilities on cross-domain tasks, we created two cross-domain datasets: Mathematics-Coding and Medical-Mathematics. For detailed experimental results, please refer to A.3. These results demonstrate that LoRA-Mixer maintains strong generalization and input-awareness capabilities on cross-domain datasets.
The impact of the RSL. We investigated the impact of our proposed RSL on the LoRA-Mixer. RSL offers three key advantages. First, it enables routing to achieve global load balancing while maintaining robust input awareness. Second, compared to auxiliary losses, RSL requires less training data. Finally, routing trained with RSL is completely independent of the original LoRAs training data.
@@
To verify the first conclusion, we conducted experiments on Medical, GSM8K, and HumanEval. As shown in Figure 4, when using RSL, the router consistently assigns higher activation weights to relevant experts, demonstrating its strong domain awareness and adaptive specialization capabilities. In contrast, when using only the auxiliary loss, the router disregards the semantics of the input and evenly distributes experts, resulting in suboptimal performance.
@@
As shown in Table 4.5, RSL achieves comparable or even superior performance using only 51.62% of the training data with auxiliary loss. We explain the suboptimal RSL results at 4k in A.16. This exceptionally high data efficiency is demonstrated by generalization bound estimation. See the Appendix A.2 for a detailed proof.
5 Conclusion.